OpenStudy (anonymous):

come up with a formula for the function f(x) whose expansion in powers of x is ((x^2)/2! )+((x^3)/3! )+((x^4)/4! )+((x^5)/5! )+...+((x^n)/n! )

OpenStudy (anonymous):

looks a lot like \(e^x\) except you are missing the first two terms

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

i would say \(f(x)=e^x-1-x\)

OpenStudy (anonymous):

it's not e^x maybe i think it's second one thankyou

OpenStudy (anonymous):

oh it is not \(e^x\) because you are missing the first two terms of the expansion

OpenStudy (anonymous):

\(e^x=1+x+\frac{x^2}{2!}+\frac{x^3}{3!}+...\) so \[e^x-1-x=\frac{x^2}{2!}+\frac{x^2}{3!}+...\]

OpenStudy (anonymous):

can you help me on another question?

OpenStudy (anonymous):

sure go ahead and post, i will help if i can

OpenStudy (anonymous):

it's two parametric function to get the volume

OpenStudy (anonymous):

probably not, but post in a new question and i bet you will get lots of replies i will try it too

OpenStudy (anonymous):

i hate integrating, it is a pain